OpenStudy (freemap):

What is the length of the segment joining the points at (1, –8) and (5, 0)? Round to the nearest tenth if necessary.

OpenStudy (freemap):

I got 11.31

OpenStudy (freemap):

@jdoe0001

OpenStudy (jdoe0001):

hmmm one sec

OpenStudy (freemap):

ok

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 1}}\quad ,&{\color{blue}{ -8}})\quad % (c,d) &({\color{red}{ 5}}\quad ,&{\color{blue}{ 0}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\) what would that give you? is not 11.31 btw

OpenStudy (wolf1728):

sq root of (16 + 64)

OpenStudy (jdoe0001):

yeap

OpenStudy (freemap):

8.94

OpenStudy (wolf1728):

That seems about right

OpenStudy (jdoe0001):

\(\bf d=\sqrt{4^2+8^2}\implies d=\sqrt{80}\implies d=\sqrt{4^2\cdot 5}\implies d=4\sqrt{5}\) or the decimal 8.94, yes

OpenStudy (freemap):

They say Round to the nearest tenth if necessary. would 9.94 still be right

OpenStudy (jdoe0001):

ohh the nearest hmm

OpenStudy (jdoe0001):

hmmm 8.94 can round up to 9.00

OpenStudy (wolf1728):

But that wouldn't be the nearest tenth

OpenStudy (jdoe0001):

nearest tenth, meaning the decimal on the right so for say 8.25 that be 8.30 but this is 8.94, so I'd think it'd make it 9.00.... since you can't make it 8.100 =)

OpenStudy (jdoe0001):

hmm wait a sec... is .94 you have to drop it down, to .90 so.. 8.90 :)

OpenStudy (wolf1728):

8.9 would be the answer

OpenStudy (freemap):

Thanks so much @jdoe0001 and @wolf1728

OpenStudy (jdoe0001):

yes, as @wolf1728 said, notice 94 is below "95" or the 5-threshold so it rounds downwards, to 90 or 8.90 or 8.9 for significant numbers